# Peeling Context from Cause for Multimodal Molecular Property Prediction

### v1

@@ 3.1 Experimental Settings
Baselines. We compare against two representative categories of methods: (i) architecture-centric models that focus on network design for molecular property prediction, including MAT (MAT) (Molecule Attention Transformer) and FP-GNN (Cai2022FPGNN) (Fingerprints and Graph Neural Networks); and (ii) causality-oriented approaches that aim to improve prediction by explicitly modeling causal structures or mitigating confounding, including DIR (DIR) (Discovering Invariant Rationales) and CAL (CAL) (Causal Attention Learning). For causality-oriented models originally designed for classification, we retain their core architectures and modify only the output head to perform regression (i.e., single-scalar prediction), preserving their causal mechanisms unchanged. See Appendix D.2 for further implementation details.
3.2 Main Results
Figure 3 summarizes the overall comparative regression performance on four molecular property datasets. CLaP consistently achieves the best MAE, MSE, and $R^{2}$ across all benchmarks. Notably, CycPeptMPDB, the most challenging benchmark due to its focus on peptide permeability, CLaP achieves a 17.0% lower MAE and 25.4% lower MSE compared to the best baseline (FP-GNN), along with an absolute $R^{2}$ improvement of 8.6 points. Empirically, this collectively highlights the strong advantage of our causal–trivial peeling under complex structure–property confounding.
Beyond CycPeptMPDB, CLaP achieves consistent improvement on ESOL, FreeSolv, and Lipo, outperforming prior architectures and causality-based models in accuracy and correlation metrics, especially on FreeSolv, where data scarcity ($\sim$600 samples) limits performance. These results further demonstrate our design’s generalizability across molecular size, data regime, and property type.
3.3 Variants and Ablations
Table 1: Ablation study on framework design across four datasets.
Setting MAE MSE $R^{2}$
ESOL
CLaP $0.4456\pm 0.0151$ $0.3583\pm 0.0095$ $0.9157\pm 0.0022$
w/o causal–trivial split $0.5094\pm 0.0528$ $0.4939\pm 0.1303$ $0.8838\pm 0.0243$
w/o correlation schedule $0.5078\pm 0.0190$ $0.4767\pm 0.0677$ $0.8879\pm 0.0159$
w/o trivial branch $0.4732\pm 0.0210$ $0.4137\pm 0.0253$ $0.9059\pm 0.0093$
FreeSolv
CLaP $0.7020\pm 0.0044$ $0.8866\pm 0.0522$ $0.9160\pm 0.0049$
w/o causal–trivial split $1.2858\pm 0.2120$ $2.5719\pm 0.8194$ $0.7562\pm 0.0777$
w/o correlation schedule $0.8779\pm 0.1003$ $1.3742\pm 0.0433$ $0.8697\pm 0.0041$
w/o trivial branch $1.0753\pm 0.0593$ $1.7338\pm 0.2021$ $0.8356\pm 0.0919$
Lipo
CLaP $0.4672\pm 0.0020$ $0.3645\pm 0.0274$ $0.7515\pm 0.0187$
w/o causal–trivial split $0.5145\pm 0.0425$ $0.4302\pm 0.0242$ $0.7240\pm 0.0062$
w/o correlation schedule $0.5562\pm 0.0195$ $0.5122\pm 0.0420$ $0.6508\pm 0.0282$
w/o trivial branch $0.5824\pm 0.0200$ $0.5348\pm 0.0115$ $0.6292\pm 0.0072$
CycPeptMPDB
CLaP $0.3056\pm 0.0080$ $0.1644\pm 0.0082$ $0.7282\pm 0.0136$
w/o causal–trivial split $0.3120\pm 0.0047$ $0.1834\pm 0.0063$ $0.6969\pm 0.0104$
w/o correlation schedule $0.3215\pm 0.0068$ $0.1866\pm 0.0087$ $0.6917\pm 0.0144$
w/o trivial branch $0.3214\pm 0.0012$ $0.2037\pm 0.0116$ $0.6634\pm 0.0192$
3.3.1 Framework design
@@
Table 1 reports ablations of core framework components. The full model integrates the depthwise correlation schedule (Eq. equation 8) and a stop–gradient trivial branch (Eq. equation 11).
@@
w/o causal–trivial split. This variant removes the correlation loss $\mathcal{L}_{\mathrm{corr}}$, the trivial loss $\mathcal{L}_{\mathrm{triv}}$, and the monotonicity penalty $\mathcal{L}_{\mathrm{mono}}$, thereby collapsing our two-branch design. Training reduces to plain regression with $\mathcal{L}_{\mathrm{pred}}$ (Eq. equation 12). Empirically, performance drops (e.g., on the CycPeptMPDB dataset, $R^{2}$ drops from 0.728 to 0.697), indicating that straightforward end-to-end fitting struggles to separate label-relevant structure and is easily distracted by spurious information.
@@
Under the batch-wise invariance view, a predictor that mixes batch-dependent context $q(x,e)$ with intrinsic signal $c(x)$ yields $\mathrm{Corr}_{e}(s,y)$ that fluctuates with $e$. Removing the split collapses the losses $\mathcal{L}_{\mathrm{corr}},\mathcal{L}_{\mathrm{triv}},\mathcal{L}_{\mathrm{mono}}$, allowing $b\!\neq\!0$ in Eq. equation 2, contradicting the invariant solution driven by Eq. equation 5. The performance drop reflects increased reliance on batch-sensitive shortcuts.
@@
w/o correlation schedule. We omit the correlation loss $\mathcal{L}_{\mathrm{corr}}$ and (Eq. equation 8), which disables the target schedule for $\rho$. The depth schedule $\{\rho_{\ell}\}$ (Eq. equation 7) and $\mathcal{L}_{\mathrm{corr}}$ (Eq. equation 8) specify a curriculum from coarse to decisive label-relevant factors.
@@
Without this schedule, the stack lacks directional guidance and drifts toward batch-specific shortcuts. It fits residual patterns with little label signal, which hurts generalization. Keeping the monotonicity term $\mathcal{L}_{\mathrm{mono}}$ is not enough because the non-decreasing constraint lacks direction and cannot steer which features to keep or discard. In the full model the schedule works as a curriculum that allocates capacity to increasingly predictive components and focuses deeper layers on label relevant variation.
@@
w/o trivial branch. We ablate the trivial branch and predict solely from the final-layer causal output $y_{\mathrm{c}}^{\star}$. This forces residual context into the causal branch and eliminates residual calibration of environment-driven variation, causing leakage of non-causal factors and a performance drop. See Appendix A.2 for a formal derivation and risk decomposition.
For instance, in a testing example with ground-truth permeability y=0.80y=-0.80, we observe $y_{\mathrm{c}}^{\star}=-0.78$ and t=0.01t=-0.01, yielding a final prediction $\widehat{y}=y_{\mathrm{c}}^{\star}+t=-0.79$. This decomposition reflects how $y_{\mathrm{c}}^{\star}$ captures the stable, mechanism-driven component, while $t$ calibrates a small, context-dependent bias. Without the trivial head, however, such residual calibration becomes unavailable. The model is forced to encode both signals into $y_{\mathrm{c}}^{\star}$, entangling causal and contextual information.
Table 2: Layers and $\rho_{\max}$ vs. performance on CycPeptMPDB.
    Setting     MAE     MSE     $R^{2}$
    Peeling Depth($L$)
    3     0.3265 $\pm$ 0.0044     0.1903 $\pm$ 0.0044     0.6854 $\pm$ 0.0073
    5     0.3056 $\pm$ 0.0080     0.1644 $\pm$ 0.0082     0.7282 $\pm$ 0.0136
    7     0.3296 $\pm$ 0.0007     0.2020 $\pm$ 0.0015     0.6662 $\pm$ 0.0025
    9     0.3297 $\pm$ 0.0070     0.1875 $\pm$ 0.0077     0.6902 $\pm$ 0.0127
    Peeling Strength($\rho_{\max}$)
    0.60     0.3131 $\pm$ 0.0075     0.1741 $\pm$ 0.0097     0.7123 $\pm$ 0.0161
    0.70     0.3162 $\pm$ 0.0040     0.1733 $\pm$ 0.0053     0.7137 $\pm$ 0.0088
    0.80     0.3056 $\pm$ 0.0080     0.1644 $\pm$ 0.0082     0.7282 $\pm$ 0.0136
    0.90     0.3133 $\pm$ 0.0034     0.1834 $\pm$ 0.0035     0.6970 $\pm$ 0.0057
    1.00     0.3136 $\pm$ 0.0183     0.1798 $\pm$ 0.0156     0.7028 $\pm$ 0.0257
3.3.2 Causal Invariance under Re-batching
Table 3: Test $R^{2}$ under $\mathsf{do}(E)$ via re-batching.
B=8B{=}8 B=16B{=}16 B=32B{=}32
$R^{2}$ (no) 0.7172 0.7282 0.7219
$R^{2}$ (shuf) 0.7120 0.7165 0.7184
Figure 4: Environment interventions via re-batching.
Setup as intervention. To evaluate the robustness of the learned representations under distributional shifts, we treat the test-time batch setup as an intervention variable. We fix the trained model and intervene on the test-time environment $E$ by changing how the same test examples are partitioned into mini-batches: $E=(B,\text{shuffle})$ with $B\!\in\!\{8,16,32\}$ and with/without shuffling. This realizes $\mathsf{do}(E=e)$ that perturbs batch boundaries and co-occurrences while holding $c(x)$ fixed. In the notation of Sec. 2.1, such interventions modify the batch-dependent context $q(x,E)$ in the local predictor $s(x,E)=ac(x)+bq(x,E)$ (Eq. equation 2) but leave the sample-intrinsic signal $c(x)$ unchanged. Our learning mechanism (depth-wise correlation targets equation 6–equation 8) is designed to drive $b\to\!0$, i.e., to make the causal readout rely on $c(x)$ and be invariant to $\mathsf{do}(E=e)$.
For each intervention we compute global correlations between $\bm{c}$} and $y$ over the test set. The correlation–depth curves (Fig. 4) remain monotonic and track the anchors for all $B$ and shuffling choices. (Table 3) Final-layer correlations lie in [0.7120, 0.7282][0.7120,\,0.7282] with total spread $\leq\!0.0162$. Since interventions change $E$ while keeping the examples fixed, this stability indicates that the learned alignment is insensitive to environment construction, as predicted by the batch-wise invariance principle supporting our claim that the predictor relies primarily on the sample-intrinsic component $c(x)$.
3.3.3 Peeling Design
We analyze how peeling depth $L$ (number of causal blocks) and strength $\rho_{\max}$ (maximum target correlation) shape performance. Pushing either to extremes yields under- or over-peeling.
Peeling depth. We vary the number of causal blocks $L\in\{3,5,7,9\}$ and observe a U-shaped response in MAE/MSE and $R^{2}$ on CycPeptMPDB in the upper block of Table 2. With too few blocks (L=3L{=}3), the hierarchy under-peels. It lacks capacity to progressively strip context, leaving causal and trivial signals entangled and hurting accuracy. With too many blocks (L=7,9L{=}7,9), the hierarchy over-peels. The longer optimization path increases variance and can misroute label-relevant patterns into the trivial stream despite the monotonicity regularizer, degrading generalization. An interior optimum at L=5L{=}5 aligns with a bias–variance view: depth must be large enough to separate signals but not so large that true causal evidence is shaved off together with context.
Peeling strength. We also examine the effect of the target-correlation parameter $\rho_{\max}\!\in\!\{0.60,0.70,0.80,0.90,1.00\}$ on CycPeptMPDB in the lower block of Table 2. Smaller targets (0.60–0.70) under-peel. The causal branch is not pushed hard enough, so contextual residue persists and errors rise. A moderate target (0.80) yields the best trade-off. Pursuing near-perfect within-batch correlation (0.90–1.00) over-peels. To satisfy an aggressive target, the model exploits batch-coupled features that should remain trivial, increasing variance and hurting generalization. Appendix A.3 provides additional evidence that an overly large target correlation causes context leakage.
3.4 Case study
Each causal block applies node-wise gating on atoms, producing $\alpha_{i}^{(\ell,m)}\!\in\![0,1]$ for atom/group $i$ in modality $m\!\in\!\{\textit{SM},\textit{PE},\textit{GE}\}$ at depth $\ell$. We take the final-layer weight as the causal score, $\pi_{i}^{(m)}=\alpha_{i}^{(L,m)}\in[0,1]$, and render causal saliency maps $\{\pi_{i}^{(m)}\}$ to inspect learned behavior.
Higher $\pi_{i}^{(m)}$ indicates atoms routed through the causal branch at the final layer. Figure 5 compares the under-peeling, optimal-peeling, and over-peeling settings. In the optimal case, saliency concentrates on chemically meaningful moieties, consistent with chemical intuition.

### v2

@@ 3.1. Experimental Setup
Baselines. We compare CLaP with two major categories of molecular property prediction methods for direct comparison: (i) architecture-centric models that emphasize network design, including representative MAT (Maziarka et al., 2019) (Molecule Attention Transformer), FP-GNN (Cai et al., 2022) (Fingerprint-enhanced GNN), GSL (Zhao et al., 2024) (Graph Structure Learning), and MolFCL (Tang et al., 2025) (fragment- and functional-group-aware contrastive learning); and (ii) causality-oriented approaches that improve prediction by modeling causal structure or mitigating confounding effects, including DIR (Wu et al., 2022) (Discovering Invariant Rationales) and CAL (Sui et al., 2022) (Causal Attention Learning). We also include CGR (Yin et al., 2025) (Causal Graph Learning), which leverages causal intervention to enhance molecular regression. For causality-based models originally proposed for classification, we preserve their causal mechanisms and adapt only the output layer for regression. See Appendix D.2 for further baseline implementation details.
As summarized in Table 2(b), the considered baselines primarily operate on single- or dual-modality inputs, focusing on 2D molecular graphs alone or augmenting them with auxiliary representations such as molecular fingerprints or 3D structures. In contrast, CLaP jointly incorporates 2D graphs, 3D structural information, and optional sequence-level HELM representations for peptides within a unified framework. From a computational perspective, the additional cost introduced by CLaP is modest. It scales linearly with the number of nodes and modalities and remains a minor overhead compared to backbone message passing in practice, since the numbers of modalities and peeling layers are small and fixed. Moreover, compared to other causality-oriented approaches that rely on manually fabricated environment interventions, CLaP avoids additional computation associated with repeated intervention-based training.
Table 2. Overview of datasets and method characteristics used in this study. (Top) Dataset statistics under in-distribution (ID) and out-of-distribution (OOD) settings. (Bottom) Comparison of computational complexity and input modalities.
(a) Dataset statistics under ID and OOD settings.
Setting Dataset Type # Samples Avg. # Atoms
ID ESOL Physical $1128$ $13.2899$
FreeSolv Physical $642$ $8.7227$
Lipo Physical $4171$ $27.0269$
CycPeptMPDB Physical $6960$ $64.8966$
OOD Half Life PK/ADME $667$ $27.6977$
Hepatocyte PK/ADME $1213$ $28.5812$
Microsome PK/ADME $1102$ $28.4519$
Solubility Physical $9982$ $17.3830$
Toxicity Safety $7385$ $16.1641$
(b) Computational complexity and input modalities across methods.
Method Computational Complexity Input Modality
FP-GNN $O(L_{back}(V+E)+F)$ 2D + FP
GSL $O(V^{2}+L_{back}(V+E))$ 2D + FP
MAT $O(L\cdot V^{2})$ 2D + 3D
CAL / CGR $O(L_{back}(V+E)+B)$ 2D
DIR $O(L_{back}(V+E)+|Env|\cdot B)$ 2D
MolFCL $O(L_{back}(V+E)+m^{2})$ 2D
CLaP $\mathbf{O(L_{back}(V+E)+L\cdot M\cdot V)}$ 2D + 3D + (HELM)
3.2. Performance Comparison
Table 1 compares CLaP against seven baselines on nine datasets spanning both ID (ESOL, FreeSolv, Lipo, CycPeptMPDB) and OOD (Half Life, Hepatocyte, Microsome, Solubility, Toxicity) settings. We report MAE and MSE. All results are averaged over 5 runs.
CLaP ranks first on 7/9 datasets in both MAE and MSE and yields substantial improvements over the strongest non–CLaP baselines. On challenging ID benchmarks such as CycPeptMPDB and FreeSolv, CLaP achieves large gains, and these improvements extend to OOD settings, including Microsome and Toxicity. On the remaining tasks, CLaP remains competitive: it attains the best MSE on Lipo, achieves the lowest MAE on Hepatocyte with near-best MSE, while Solubility is the only dataset where MolFCL leads on both metrics.
These results indicate that causal–context peeling—via the correlation schedule and residual context calibration—consistently improves data efficiency and robustness across both ID and OOD regimes. This empirically supports the effectiveness of the proposed framework in disentangling and effectively leveraging causal and contextual information for molecular property prediction.
3.3. Variants and Ablations
3.3.1. Framework design
@@
Table 1 also reports ablations of framework components. The framework integrates depthwise correlation schedule (Eq. equation 5) and the context branch (Eq. equation 8).
@@
w/o causal–context split. This variant removes the correlation loss $\mathcal{L}_{\mathrm{corr}}$, the context loss $\mathcal{L}_{\mathrm{ctx}}$, and the monotonicity penalty $\mathcal{L}_{\mathrm{mono}}$, thereby collapsing our two-branch design. Training reduces to plain regression with $\mathcal{L}_{\mathrm{pred}}$ (Eq. equation 9). Empirically, performance drops (e.g., on the CycPeptMPDB dataset), indicating that straightforward end-to-end fitting struggles to separate label-relevant structure and is easily distracted by spurious information.
@@
Under the batch-wise invariance view, a predictor that mixes batch-dependent context $q(x,e)$ with intrinsic signal $c(x)$ yields $\mathrm{Corr}_{e}(s,y)$ that fluctuates with $e$. Removing the split collapses the losses $\mathcal{L}_{\mathrm{corr}},\mathcal{L}_{\mathrm{ctx}},\mathcal{L}_{\mathrm{mono}}$, allowing $b\!\neq\!0$ in Eq. equation 1, contradicting invariant solution driven by Eq. equation 3. The performance drop reflects increased reliance on batch-sensitive shortcuts.
@@
w/o correlation schedule. We omit the correlation loss $\mathcal{L}_{\mathrm{corr}}$ and (Eq. equation 5), which disables the target schedule for $\rho$. The depth schedule $\{\rho_{\ell}\}$ and $\mathcal{L}_{\mathrm{corr}}$ (Eq. equation 5) specify a curriculum from coarse to decisive label-relevant factors.
@@
Without this schedule, the stack lacks directional guidance and drifts toward batch-specific shortcuts. It fits context patterns with the label signal, which hurts generalization. Keeping the monotonicity term $\mathcal{L}_{\mathrm{mono}}$ is not enough because the non-decreasing constraint lacks direction and cannot steer which features to keep or discard. In the full model the schedule works as a curriculum that allocates capacity to increasingly predictive components and focuses deeper layers on label-relevant variation.
@@
w/o context branch. We ablate the context branch and predict solely from the final-layer causal output $y_{\mathrm{c}}^{\star}$. This eliminates residual calibration of environment-driven variation into prediction. See Appendix A.2 for a formal derivation and risk decomposition.
To illustrate, consider a molecule whose aqueous solubility is primarily driven by a polar functional group (e.g., a nitro or hydroxyl group), while the overall scaffold or ring system introduces secondary, context-dependent effects. In this case, the causal prediction $y_{\mathrm{c}}^{\star}$ captures the contribution of the functional group but may deviate slightly from the ground-truth label $y$ due to scaffold-dependent or batch-specific factors. The context branch learns to produce a residual term $y_{\text{ctx}}\approx y-y_{\mathrm{c}}^{\star}$, calibrating the final prediction as $\widehat{y}=y_{\mathrm{c}}^{\star}+y_{\text{ctx}}$. Without this branch, chemical information related to scaffold- and batch-dependent effects cannot be incorporated to correct deviations in $y_{\mathrm{c}}^{\star}$, reducing overall predictive fidelity.
Figure 3. Effect of causal peeling across modalities and consistency weight on prediction performance. Lower is better.
3.3.2. Multi-modality Design
In this section, we first examine how the proposed causal–context split behaves under different modality configurations. We consider both unimodal and multimodal settings, where CycPeptMPDB provides three complementary views per sample(SM, PE, GE), while ESOL, FreeSolv, and Lipo each provide two views (SM and GE). Fig. 3 compares unimodal variants with multimodal fusion models, with and without the causal split. Across datasets with two or more modalities, enabling the causal–context split consistently improves both accuracy and stability over unimodal baselines. These results indicate that while multimodality supplies richer input signals, the causal–context split plays a role in structuring and exploiting multimodal information, allowing CLaP to leverage different modalities in a complementary manner.
(a) Causal analysis across molecular scaffolds.
(b) Environment interventions via re-batching.
Figure 4. Causal evaluation under scaffold and environment shifts.
Next, we investigate whether explicitly enforcing cross-modal agreement is beneficial beyond the proposed causal split. To this end, we introduce a cosine-similarity penalty between modality-specific embeddings, weighted by $\lambda_{\text{cons}}$, and evaluate its effect (Fig. 3). Across datasets, adding this penalty degrades performance compared to the default setting without cross-modal consistency ($\lambda_{\text{cons}}=0$), with larger values of $\lambda_{\text{cons}}$ leading to larger drops.
Overall, the proposed split improves how information from multiple modalities is utilized effectively. By separating label-relevant signals from contextual residue, the model can more selectively leverage complementary cues across modalities. In contrast, without this separation, multimodal signals remain entangled, leading to weaker performance despite access to multiple inputs.
3.4. Causal Evaluation under Shifts
To examine whether CLaP exhibits intended causal behavior beyond predictive accuracy, we conduct evaluations under controlled environment shifts and analyze how the causal and context branches respond to changes in the environment at test time.
3.4.1. Causal Effects across Structural Environments
Evaluation setup.
To evaluate whether the proposed CLaP can separate environment-level structure from sample-specific variation, we analyze its behavior from a structural hierarchy perspective. The goal of this experiment is to examine whether the model can effectively disentangle shared scaffold-level effects from residual molecule-level variation across different structural scales.
We consider prediction results for all molecular properties in the test set. For analytical clarity, we conceptually distinguish two sources of structure-related variation. The first corresponds to scaffold structures defined by the Bemis–Murcko framework, which represent stable and shared environmental factors within each structural cluster. The second consists of local structural variations attached to the scaffold, such as functional groups or substitution patterns, whose effects are typically expressed as molecule-specific differences. This partition is introduced for causal analysis rather than as an assumption about underlying chemical mechanisms.
For each property, molecules are grouped into clusters according to their scaffolds, and only clusters containing at least five molecules are retained to ensure statistical stability. Within each cluster, we compute the cluster-level means of the causal branch output, the context branch output, and the ground-truth property values. This cluster-wise averaging operation statistically marginalizes molecule-level local variation, thereby isolating scaffold-level shared structural information in expectation across clusters.
To assess whether scaffold-induced systematic variation is correctly allocated across branches, we examine the correlations between cluster-level branch outputs and the aggregated labels.
Observed behavior.
As shown in Fig. 4(a), scaffold-induced baseline shifts are captured by the context branch, which shows strong correlation with the labels, whereas the causal branch remains almost invariant across scaffolds and exhibits weaker correlation.
This pattern arises from the scaffold-level aggregation used in this evaluation and provides direct evidence for the intended causal separation in our model. By averaging predictions and labels within each scaffold cluster, molecule-specific variations—primarily functional-group-level effects—are statistically marginalized in expectation, leaving scaffold-level shared variation as the dominant signal across clusters. Since scaffold identity represents a stable environmental factor shared by all molecules in a cluster, its effect manifests as systematic shifts in the cluster-level mean label and is therefore captured by the context branch. In contrast, the causal branch encodes sample-specific residual variation and becomes approximately invariant after within-cluster averaging, resulting in its diminished correlation with the aggregated labels.
Notably, for certain scaffold-dominated properties such as solubility, the causal structure can be inherently multi-scale. Beyond functional-group-level polarity, molecule-specific structural factors related to scaffold geometry—such as steric effects or substitution patterns—may introduce residual variation within scaffold clusters. The causal branch may retain weak but non-negligible correlation in these cases, reflecting remaining sample-intrinsic effects in practice rather than leakage of scaffold-level shared information.
These results indicate that, under scaffold-level marginalization, CLaP correctly allocates environment-level structural effects to the context branch while preserving molecule-specific variation in the causal branch in this analysis as intended by design.
3.4.2. Causal Invariance under Re-batching
Evaluation setup.
To test whether the learned causal representation exhibits invariance under interventions on the environment, we conduct a controlled evaluation by manipulating the test-time batching scheme. The batch construction is treated explicitly as an environment variable $E$. Concretely, we intervene on $E$ by altering how the same test examples are grouped into mini-batches, with $E=(B,\text{shuffle})$, where $B\!\in\!\{8,16,32\}$ and shuffling is applied or not. This realizes a family of interventions $\mathsf{do}(E=e)$ that perturb batch boundaries and co-occurrence statistics, while preserving the underlying sample-intrinsic signal $c(x)$ by construction.
Under the batch-wise invariance formulation in Sec. 2.1, such interventions affect only the batch-dependent context component $q(x,E)$ in the local predictor $s(x,E)=ac(x)+bq(x,E)$ (Eq. equation 1), while leaving $c(x)$ unchanged. The depth-wise correlation schedule in our training objective (Eqs. equation 4–equation 5) is designed to drive $b\rightarrow 0$, encouraging the causal branch to depend on $c(x)$ and remain approximately invariant under $\mathsf{do}(E=e)$.
Observed behavior.
For each intervention, we compute the global Pearson correlation between the causal prediction $y_{\mathrm{c}}^{\star}$ and the ground-truth labels over the entire test set. As shown in Fig. 4(b), the resulting correlation–depth curves remain tightly aligned across all batch sizes and shuffling configurations in practice.
At the final layer, correlations concentrate within a narrow range, indicating that the causal branch is insensitive to how test-time environments are constructed. Minor deviations from the anchor correlation can be attributed to context-induced effects, such as scaffold composition or other batch-level statistics. Importantly, these effects are explicitly modeled and absorbed by the context branch, allowing the causal path to preserve an invariant predictive signal across interventions. Since the interventions modify the environment variable $E$ while keeping the underlying examples fixed, the observed stability provides direct evidence that the learned causal representation satisfies the desired invariance property.
Together with the structural environment analysis, these two experiments provide complementary evidence of causal behavior. The structural analysis verifies that environment effects are allocated to the context branch, while the re-batching intervention demonstrates that the causal branch remains invariant under environment shifts, confirming causal–context disentanglement in CLaP.